# Exploiting Minority Pseudo-Labels for Semi-Supervised Semantic Segmentation in Autonomous Driving

### v2

@@ 4.3 Analysis
Table 4: Improvement of each proposed module: $\mathcal{L}_{s}$ represents the supervised loss, while $\mathcal{L}_{u}$ denotes the unsupervised loss. DM signifies a dual mean teacher architecture, PS denotes the pixel selection strategy, and DCL indicates dual contrastive learning. The results are obtained using 1/8 labeled data from both the Cityscapes and PASCAL VOC 2012 datasets, employing DeepLabV3+ with ResNet50.
@@
$\mathcal{L}_{s}$ $\mathcal{L}_{u}$ DM PS DCL Cityscapes
@@
Ablation Studies. Table 4 presents the contributions of each module within our framework. Each module contributes significantly to enhancing semi-supervised semantic segmentation performance. Initially, the baseline $\mathcal{L}_{s}$, which is trained solely on labeled data, achieves a certain mIoU score. Then, we adopt the traditional mean-teacher architecture to introduce unlabeled data $\mathcal{L}_{u}$ and improve the performance of the model. However, the introduction of the dual mean teacher architecture (DM) individually boosts the mIoU to 73.37% for Cityscapes and 73.92% for PASCAL VOC 2012, indicating improvements of 6.63% and 7.09%, respectively.
Furthermore, incorporating the pixel selection strategy (PS) and dual contrastive learning (DCL) modules leads to additional performance gains. The PS and DCL modules surpass the baseline by 7.22% and 7.45%, respectively, for Cityscapes and by 8.04% and 7.43%, respectively, for PASCAL VOC 2012.
Upon integrating these modules, the overall performance escalates, reaching 75.14% for Cityscapes and 75.49% for PASCAL VOC 2012. This comprehensive analysis underscores the effectiveness of the dual mean teacher architecture, pixel selection strategy, and dual contrastive learning in enhancing semi-supervised semantic segmentation.
Effect of Pixel Selection Strategy. As shown in Table 5, when all the settings are the same, selecting consistent and highly mismatched pseudo-labels yields the best performance. This finding indicates that highly mismatched pseudo-labels have a high probability of being correct once a prediction is made due to their difficulty in learning. In addition, because the two modules are different, if they are consistent on a pixel, there is a high probability that they will be correct.
@@
Table 5: Effect of pixel selection strategy. $\hat{y}^{Cons}_{u}$ indicates pseudo-labels with consistency. $\hat{y}^{Lmis}_{u}$ indicates low mismatched pseudo-labels. $\hat{y}^{Hmis}_{u}$ indicates highly mismatched pseudo-labels. The results are derived from the Cityscapes dataset using 1/8 of the labeled data, employing DeepLabV3+ with ResNet50.
$\hat{y}^{Cons}_{u}$ $\hat{y}^{Lmis}_{u}$ $\hat{y}^{Hmis}_{u}$ mIoU(%)
71.43
72.61
75.14
73.56
@@
Effect of the Memory Bank Size $N$. As illustrated in Fig. 6(a), the performance of the model is not significantly affected by variations in the size of the memory bank when other settings are kept constant. Minor adjustments to the memory bank size can result in slight changes in model performance. However, using an excessively large memory bank, such as a size of 512, may degrade performance. It’s important to note that setting $N=0$ indicates the removal of contrastive learning from the framework. Since all features in the memory bank are utilized during the training, the computational load is directly proportional to the size $N$. To balance performance and computational efficiency, we have set the default memory bank size to 256.
Effect of Unsupervised Loss Weight $\lambda_{u}$. In Fig. 6(b), we explore the effect of the unsupervised loss weight $\lambda_{u}$ on the final segmentation performance. A weight of 0 indicates that the unlabeled data do not participate in training. The model is not sensitive to the unsupervised loss weight in a specific range and shows fluctuations while the overall performance is improved. The initial experiment involves setting the weight to 0.5, which resulted in an mIoU of 72.5%. Subsequently, we increase the weight to 1.0, placing equal emphasis on the unsupervised loss weight relative to the supervised loss weight. This adjustment leads to an improvement in the mIoU, achieving a value of 75.14%. Further exploration involves increasing the weight to 1.5, this adjustment yields a slightly lower mIoU of 73.94%.
@@
Figure 6: Effect of hyper-parameters. (a) Effect of the memory bank size $N$; (b) Effect of unsupervised loss weight $\lambda_{u}$. The results are derived from the Cityscapes dataset using 1/8 of the labeled data, employing DeepLabV3+ with ResNet50.
@@
Figure 8: Qualitative results of STPG framework using 1/8 labeled data from the PASCAL VOC 2012 training set. (a) Images from the validation set. (b) The results of training exclusively with labeled data. (c) The results achieved with STPG. (d) Ground-truths. All experiments are conducted using DeepLabV3+ with ResNet50.
4.4 Qualitative Results
Fig. 7 and Fig. 8 illustrate the qualitative results of STPG on the Cityscapes and PASCAL VOC 2012 datasets, respectively. Notably, our method shows significant improvements over training with labeled data only, enhancing the performance of semi-supervised learning.
5 Conclusion
In this paper, we propose a new framework for semi-supervised semantic segmentation for autonomous driving. The STPG framework with professional and general training modules can reduce model coupling and error accumulation and retain more semantic information for minority classes. Within the framework, a dual contrastive learning can alleviate the influence of majority class dominance in the feature space. Our approach is easily adaptable to various semi-supervised semantic segmentation networks. The experimental results show that STPG surpasses the state-of-the-art techniques.
Nevertheless, as with other semi-supervised learning approaches, the training process of our method is more time-consuming than that of fully supervised methods, and the correctness of pseudo-labels can greatly affect the effectiveness of contrastive learning, especially when features from different classes are similar in the feature space. Resolutions to these limitations can be explored in future studies.
Declarations
Authors contribution statement Yuting Hong: Conceptualization, Methodology, Software, Writing - Original Draft. Hui Xiao: Investigation. Huazheng Hao: Investigation. Xiaojie Qiu: Software, Writing - Review & Editing. Baochen Yao: Methodology, Software. Chengbin Peng: Conceptualization, Investigation, Methodology, Project Administration, Software, Writing - Original Draft, Funding acquisition, Resources, Writing - Review & Editing, Supervision.
Competing interests All the authors declare that they have no competing financial interests or personal relationships that could influence the work reported in this paper.
Ethical approval and consent to participate The work follows appropriate ethical standards in conducting research and writing the manuscript. This work presents computational models trained with publicly available data, for which no ethical approval was required.
Data availability and access The data used to support the findings of this study is available from the corresponding author upon request.

### v3

@@ IV-C Analysis
TABLE IV: Improvement of each proposed module: $\mathcal{L}_{s}$ represents the supervised loss, while $\mathcal{L}_{u}$ denotes the unsupervised loss. DM signifies a dual mean teacher architecture, PS denotes the pixel selection strategy, and ACL indicates anchor contrastive learning. The results are obtained using 1/8 labeled data from both the Cityscapes and PASCAL VOC 2012 datasets, employing DeepLabV3+ with ResNet50.
@@
$\mathcal{L}_{s}$ $\mathcal{L}_{u}$ DM PS ACL Cityscapes
@@
Ablation Studies. Table IV presents the contributions of each module within our framework. Each module contributes significantly to enhancing semi-supervised semantic segmentation performance. Initially, the baseline $\mathcal{L}_{s}$, which is trained solely on labeled data, achieves a certain mIoU score. Then, we adopt the traditional mean-teacher architecture to introduce unlabeled data $\mathcal{L}_{u}$ and improve the performance of the model. However, the introduction of the dual mean teacher architecture (DM) individually boosts the mIoU to 73.37% for Cityscapes and 73.92% for PASCAL VOC 2012, indicating improvements of 6.63% and 7.09%, respectively. Furthermore, incorporating the pixel selection strategy (PS) and anchor contrastive learning (ACL) modules leads to additional performance gains. The PS and ACL modules surpass the baseline by 7.22% and 7.45%, respectively, for Cityscapes and by 8.04% and 7.43%, respectively, for PASCAL VOC 2012. Upon integrating these modules, the overall performance escalates, reaching 75.14% for Cityscapes and 75.49% for PASCAL VOC 2012. This comprehensive analysis underscores the effectiveness of the dual mean teacher architecture, pixel selection strategy, and anchor contrastive learning in enhancing semi-supervised semantic segmentation.
Effect of Pixel Selection Strategy. As shown in Table V, our ablation study highlights the critical role of pseudo-label selection. A baseline using only high-quality, consistently predicted pseudo-labels establishes a solid foundation but achieves a modest mIoU of 71.43%, as detailed in Row (a). This suggests that relying exclusively on easy examples limits model generalization. The key to our improved performance lies in strategically leveraging uncertainty. By incorporating high-mismatch pseudo-labels, which are identified at the class level, alongside the high-quality ones, performance significantly increases to a peak of 75.14% mIoU in Row (b). Adding high-mismatch pseudo-labels consistently boosts performance, improving mIoU by +3.71% in Row (b) and +0.95% in Row (d). This result validates our central hypothesis: uncertain, boundary-defining samples are more valuable for model improvement than deterministic ones alone. We posit that these high-mismatch pseudo labels, often generated for hard-to-classify minority classes, are rich in valuable boundary information. This is because these classes are inherently prone to being misclassified as more dominant, over-confident classes, thus yielding higher mismatch scores. Conversely, other configurations prove less effective. While adding low-mismatch labels shows a slight improvement over the baseline in Row (c), combining both low- and high-mismatch labels in Row (d) leads to a performance drop relative to our best result. This demonstrates that it is specifically the high-mismatch samples that drive the improvement.
@@
TABLE V: Effect of pixel selection strategy. $\hat{y}^{Hq}_{u}$ indicates high-quality pseudo-labels. $\hat{y}^{Lmis}_{u}$ indicates low-mismatch pseudo-labels. $\hat{y}^{Hmis}_{u}$ indicates high-mismatch pseudo-labels. The results are derived from the Cityscapes dataset using 1/8 of the labeled data, employing DeepLabV3+ with ResNet50.
$\hat{y}^{Hq}_{u}$ $\hat{y}^{Lmis}_{u}$ $\hat{y}^{Hmis}_{u}$ $mIoU(\%)$ $\Delta_{mIoU}(\%)$
(a) 71.43
(b) 75.14 +3.71
(c) 72.61
(d) 73.56 +0.95
Model Decoupling Analysis. To rigorously characterize the divergence between the student and teacher models, we define the mismatch rate $\mathrm{MR}$ as the proportion of evaluation samples $x_{i}$ whose classification predictions differ between the two models:
$\mathrm{MR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\text{argmax}f_{\theta}(x_{i})\neq\text{argmax}f_{\xi}(x_{i})),$ (20)
where $N$ is the total number of evaluate samples.
As shown in Fig. 9, during training, we periodically measured the mismatch rate between the student and teacher models for both STPG and FixMatch, taking measurements every 4 epochs. The generated plot shows that STPG maintains a consistently higher mismatch rate, revealing a looser student–teacher coupling than that of FixMatch. This higher mismatch rate is consistent with STPG’s design, where cross-teacher supervision and the complementary objectives of the two students reduce parameter coupling and encourage more diverse feature representations.
Figure 9: Mismatch Rate comparison between STPG and FixMatch. The mismatch rate between the student and teacher models was measured every 4 epochs during training for both methods. The plot indicates that STPG consistently exhibits a higher mismatch rate, suggesting a weaker coupling between its student and teacher models compared to FixMatch.
@@
Sensitivity of the Memory Bank Size $N$. As illustrated in Fig. 10(a), the performance of the model is not significantly affected by variations in the size of the memory bank when other settings are kept constant. Minor adjustments to the memory bank size can result in slight changes in model performance. However, using an excessively large memory bank, such as a size of 512, may degrade performance. Since all features in the memory bank are utilized during the training, the computational load is directly proportional to the size $N$. To balance performance and computational efficiency, we have set the default memory bank size to 256.
Sensitivity of Prototype Momentum Coefficient $\alpha$. Fig. 10(b) presents the sensitivity of the model to the prototype momentum coefficient $\alpha$. The results show that the model maintains stable performance across a broad range of $\alpha$ values. When $\alpha$ is between 0.9 and 0.999, the mIoU remains consistently high, varying from 74.49% to 75.14%, with the best score obtained at $\alpha=0.99$. Lower values such as $\alpha=0.5$ and $\alpha=0.8$ lead to slight decreases in mIoU, yielding 74.21% and 74.29% respectively. These observations indicate that the method is generally robust to the choice of $\alpha$, while a moderately large momentum around 0.99 provides the most reliable performance.
Sensitivity of Loss Weight $\lambda_{u}$ and $\lambda_{ctr}$. We comprehensively explore the impact of both the unsupervised loss weight $\lambda_{u}$ and the contrastive loss weight $\lambda_{ctr}$ on the final segmentation performance. The results are presented in Fig. 10(c) and Fig. 10(d) respectively. For $\lambda_{u}$, when $\lambda_{u}$ is set to 0.5, reaching an mIoU of 74.51%. The optimal performance occurs at $\lambda_{u}=1.0$, where the mIoU peaks at 75.14%. This suggests that an equal emphasis on unsupervised and supervised losses is most beneficial within this experimental setup. However, further increasing $\lambda_{u}$ to 1.5 leads to a slight decrease in performance to 74.94%, indicating that excessive weighting of unsupervised loss can eventually introduce diminishing returns. Regarding $\lambda_{ctr}$, introducing a small weight of $\lambda_{ctr}=0.1$ significantly boosts the performance to 75.14%, matching the highest mIoU observed. As $\lambda_{ctr}$ is further increased to 0.5, the mIoU remains robust at 75.06%. Even at $\lambda_{ctr}=1.0$, it holds strong at 74.89%. This demonstrates that contrastive learning consistently contributes positively to segmentation performance across a range of weights, with $\lambda_{ctr}=0.1$ appearing to be the sweet spot for maximum benefit. In summary, both unsupervised learning and contrastive learning are crucial components for enhancing segmentation performance. The model shows a clear sensitivity to $\lambda_{u}$, with an optimal point empirically found at 1.0. Conversely, the model is relatively robust to the choice of $\lambda_{ctr}$within the tested range, indicating that even a modest incorporation of contrastive loss provides significant and consistent benefits.
@@
Figure 10: Sensitivity of hyper-parameters. (a) Sensitivity of the memory bank size NN, (b)Sensitivity of Prototype Momentum Coefficient $\alpha$. (c) Sensitivity of unsupervised loss weight $\lambda_{u}$, (d) Sensitivity of Contrastive Loss Weight $\lambda_{ctr}$. The results are derived from the Cityscapes dataset using 1/8 of the labeled data, employing DeepLabV3+ with ResNet50.
@@
Figure 12: Qualitative results of STPG framework using 1/8 labeled data from the CamVid training set. (a) Images from the validation set. (b) The results of training exclusively with labeled data. (c) The results achieved with STPG. (d) Ground-truths. All experiments are conducted using DeepLabV3+ with ResNet50.
IV-D Qualitative Results
Fig. 11 and Fig. 12 present the qualitative results of STPG on the Cityscapes and CamVid datasets, respectively. It is evident that our method achieves substantial improvements over training using only labeled data, particularly in semi-supervised settings. Notably, significant performance gains are observed for minority classes, demonstrating the effectiveness of STPG in handling class imbalance and enhancing overall segmentation accuracy.
V Conclusion
In this work, we propose a novel framework for semi-supervised semantic segmentation to enhance fine-grained road scene understanding in intelligent transportation systems, especially when labeled data is limited. The framework incorporates both specialized and general training modules, which help reduce model coupling and error accumulation while preserving richer semantic information for minority classes. Furthermore, the introduction of anchor contrastive learning alleviates the dominance of majority classes in the feature space. In intelligent transportation systems, accurate perception of all types of road users, such as pedestrians, cyclists, and traffic signs, is essential to ensure the safety and effectiveness of autonomous driving. Our method significantly improves the recognition performance of minority classes without sacrificing the accuracy of majority classes. This balanced and dependable performance across all categories makes the proposed approach well-suited for real-world scenarios.